### v1

# The Unreasonable Effectiveness of Easy Training Data for Hard Tasks

Abstract
How can we train models to perform well on hard test data when hard training data is by definition difficult to label correctly? This question has been termed the scalable oversight problem and has drawn increasing attention as language models have continually improved. In this paper, we present the surprising conclusion that current language models often generalize relatively well from easy to hard data, even performing as well as “oracle” models trained on hard data. We demonstrate this kind of easy-to-hard generalization using simple training methods like in-context learning, linear classifier heads, and QLoRA for seven different measures of datapoint hardness, including six empirically diverse human hardness measures (like grade level) and one model-based measure (loss-based). Furthermore, we show that even if one cares most about model performance on hard data, it can be better to collect and train on easy data rather than hard data, since hard data is generally noisier and costlier to collect. Our experiments use open models up to 70b in size and four publicly available question-answering datasets with questions ranging in difficulty from 3rd grade science questions to college level STEM questions and general-knowledge trivia. We conclude that easy-to-hard generalization in LMs is surprisingly strong for the tasks studied, suggesting the scalable oversight problem may be easier than previously thought. [^1] Our code is publicly available at: https://github.com/allenai/easy-to-hard-generalization.
1 Introduction
Figure 1: A model trained on easy data (e.g., 3rd Grade problems) does almost as well on hard tasks (College problems) as a model trained on those hard tasks (the College bar). Results shown for Mixtral-8x7B prompted with $k=10$ examples, averaged over 5 random seeds.
It is difficult to supervise LMs (i.e., train LMs to give correct outputs) in specialized domains of human knowledge, because it is difficult to correctly label data in such domains. Labeling difficulty manifests itself in both time to annotate (and thus cost) and label noise Lease (2011); Northcutt et al. (2021). Labeling difficulty becomes severe when specific expertise is required Sambasivan et al. (2021). For example, for sufficiently specific physics problems, PhD holders and PhD students can make errors on as many as 40% of (objective) problems Rein et al. (2023). As more NLP benchmarks focus on challenging domain-specific tasks, having access to large human-labeled training corpora for these tasks may become increasingly infeasible (e.g., existing benchmarks like MMLU Hendrycks et al. (2020) and GPQA Rein et al. (2023) do not come with training data). The question arises: how can we train models to solve hard problems when correctly labeling enough hard data for training is difficult? This problem is an example of the scalable oversight problem, which concerns how to give a good reward signal to a model when it is difficult to assess if its outputs are correct Amodei et al. (2016).
In this paper, we study the problem of easy-to-hard generalization, which is relevant for determining how challenging the scalable oversight problem is in the first place. Easy-to-hard generalization refers to model performance on hard test data when trained only on easy training data, defined according to some human hardness measure (like grade level). Since gathering training data in domains like graduate level STEM fields is expensive and time-consuming, it would clearly be useful if we could improve model performance in these domains by only training models on cleanly labeled data from simpler domains, like high school STEM questions. To assess how well current LMs generalize from easy to hard data, we train [^2] By “training” models, we refer to fitting models to data via in-context learning (ICL), parameter efficient finetuning (QLoRA), or by training a linear classifier head. models on easy data and test them on hard data (“easy-to-hard”), then compare them to an “oracle” upper bound and unsupervised lower bound. The “oracle” upper bound is a model that has access to labeled hard data for training (“hard-to-hard”), while the unsupervised lower bound is a model that is prompted zero-shot to answer questions (“unsupervised-to-hard”). The final metric we are interested in is the Supervision Gap Recovered (SGR):
$\frac{\textrm{Easy}-\textrm{Unsupervised}}{\textrm{Hard}-\textrm{Unsupervised}}$
where Easy, Hard, and Unsupervised refer to model performance on hard test data when trained on easy data, hard data, or no data (zero-shot), respectively. This metric takes a value of 100% when training on easy data is as effective as training on hard data, and it is 0% when a model trained on easy data is no better than prompting a model zero-shot.
Figure 2: Supervising Llama-2-70b with easy data (left: high school level computer science problems; right: math problems with 2-3 reasoning steps) can enable generalization to hard data (left: a college level computer science problem; right: a math problem involving 8 reasoning steps).
Our main result is that models generalize surprisingly well from easy to hard data, often performing almost as well as a model trained on hard data (illustrated in Fig. 1). In experiments with ARC Clark et al. (2018), MMLU Hendrycks et al. (2020), GSM8k Cobbe et al. (2021), and StrategyQA Geva et al. (2021), we find that the Supervision Gap Recovered is usually between 70% and 100%, meaning that easy supervision is at least 70% as good as hard supervision for hard test performance (see Fig. 2 for example problems). These results are robust across (1) model scale between 7b and 70b parameters, (2) six different human hardness measures and a model-based measure, (3) four datasets/tasks, and (4) several training methods including in-context learning with and without chain-of-thought reasoning Brown et al. (2020); Wei et al. (2022), QLoRA Dettmers et al. (2023), and linear classifier heads Belinkov (2022). Overall, our results suggest that current LMs generalize relatively well to held-out data across human difficulty levels even when trained on data that is measurably easier than the test data.
This finding implies that easy supervision may outperform hard supervision when (1) within some data collection budget, a greater quantity of easy data can be collected than hard data, or (2) easy data can be labeled with lower error rates than hard data (see Sec. 5.3). In these settings, we are able to train models that perform better on hard test data when trained on easy data rather than hard data.
Lastly, we aim to understand how easy-to-hard generalization changes with model scale and the gap between train and test hardness (Sec. 5.4). We find that (1) the Supervision Gap Recovered is highly robust across model scales between 7b and 70b parameters, and (2) easy-to-hard performance may begin to decline when the gap between train and test hardness becomes sufficiently large.
The remainder of the paper is organized along the following research questions:
  1. RQ1:
    How Can We Measure Data Hardness? Do Different Approaches Agree?
  2. RQ2:
    Can We Do Well on Hard Data by Training on Easy Data?
  3. RQ3:
    What Are the Cost-Benefit Tradeoffs of Collecting Easy vs. Hard Training Data?
  4. RQ4:
    Is Easy-To-Hard Generalization Consistent Across Model Scale and Train-Test Hardness Gap Size?
We summarize our main conclusions below:
  1. 1.
    Our six human hardness measures and one model-based measure are empirically diverse, and model performance declines on harder test data for each measure.
  2. 2.
    LMs generalize surprisingly well from easy-to-hard data, closing 70%-100% of the performance gap between unsupervised-to-hard and hard-to-hard accuracy.
  3. 3.
    We show that it is often better to train on easy data when hard data is more expensive to collect or has noisier labels.
  4. 4.
    The Supervision Gap Recovered is highly robust across model scale. Easy-to-hard performance may begin to decline when the train-test hardness gap is large enough.
2 Related Work
Curriculum Learning. Curriculum learning has historically concerned itself with model performance on hard data points Bengio et al. (2009). Previous work in this area has argued that learning from easy data first is helpful for models later learning more complex concepts and therefore performing better on hard data Xu et al. (2020). Yet, other studies also find that a hard-then-easy ordering can work well Maharana and Bansal (2022), as can non-monotonic difficulty orderings Elgaar and Amiri (2023). Whereas curriculum learning aims to improve hard test performance by optimally ordering the training data, we simply aim to investigate how well models generalize to hard data based on the hardness of the training data. Our results suggest that pretrained LMs generalize surprisingly well from easy to hard data, potentially alleviating the need for heavily engineered training curricula.
Compositional Generalization. Work in compositional generalization has previously shown that neural networks struggle to generalize to problems that require combining reasoning steps in ways not seen exactly during training Lake and Baroni (2018); Bogin et al. (2022); Zhou et al. (2023b). Further work has begun characterizing the conditions under which models generalize to compositionally more difficult problems. For instance, Transformers will generalize better on classes of algorithmic problems whose solutions can be written in RASP, meaning the programs can be implemented exactly by a Transformer forward pass Zhou et al. (2023a). Recurrent test-time computation also appears to be quite valuable for generalizing to problems requiring more reasoning steps than those seen during training Schwarzschild et al. (2021); Bansal et al. (2022). However, even GPT-3.5 with Chain-of-Thought prompting can struggle to generalize to simple mathematical problems requiring more reasoning steps than seen during training Dziri et al. (2023). Our results are not inconsistent with these studies, but instead demonstrate that relative to an unsupervised-to-hard lower bound and hard-to-hard upper bound, easy-to-hard performance on compositional reasoning problems is often surprisingly good (Sec. 5.2). In Sec. 5.1, we demonstrate the more basic conclusion (agreeing with past studies) that models perform worse on test data requiring a higher number of reasoning steps.
Easy-to-Hard Generalization. Amodei et al. (2016) motivate the scalable oversight problem by pointing out how it could be challenging to give a proper reward signal to a model when it is difficult to assess if its outputs are correct. Assessing easy-to-hard generalization provides useful context for understanding the difficulty of the scalable oversight problem, as it tells us how we can expect models to generalize from a setting where we can properly supervise them to one where we cannot. Past work evaluates easy-to-hard generalization in NLP using model-based hardness measures Swayamdipta et al. (2020) and number of compositional reasoning steps required to solve a problem Fu et al. (2022). Swayamdipta et al. (2020) show that BERT models perform worse on commonsense reasoning tasks when trained on easy data rather than hard data according to a loss-based metric resembling minimum description length Perez et al. (2021). Fu et al. (2022) show a similar result with GPT3 models for StrategyQA and GSM8k, finding that prompting with “complex” examples does better than “simple” examples, where examples are categorized according to the number of reasoning steps in the annotated human reasoning chain for a problem. Relative to these works, we study easy-to-hard generalization with (1) a greater number of human hardness measures, including grade level, expert rating, required cognitive skills, question length, answer length, and number of reasoning steps, as well as a model-based metric, (2) multiple datasets involving science question answering, compositional reasoning, and mathematical reasoning, and (3) multiple model sizes for understanding scaling trends. In contrast to these works, we show that in a number of settings easy-to-hard generalization is comparable to hard-to-hard generalization.
In concurrent work, Burns et al. (2023) present results on a related “weak-to-strong” generalization problem, with similar motivation to our work. They also present results specifically on easy-to-hard generalization for NLP tasks using a model-based hardness measure. In contrast to Burns et al. (2023), (1) we define our main performance metric (Supervised Gap Recovered) using an unsupervised model as the baseline performance rather than a weaker model as the baseline performance, which is important when an unsupervised stronger model will greatly outperform a supervised weaker model (as is observed in our experiments); (2) we use human hardness measures in addition to model-based hardness, which is important when the two may not correlate strongly (see our Fig. 3); and (3) we use publicly available datasets and open-source models rather than unidentified “NLP tasks” and API-gated models.
ARC MMLU-STEM-5 StrategyQA GSM8k
$n=4521$ $n=1746$ $n=2290$ $n=8792$
Grade Level (3-8) Grade Level (HS vs. College) Grade Level Grade Level
Difficulty Score (1-3) Difficulty Score Difficulty Score Difficulty Score
Bloom Skill (1-5) Bloom Skill Bloom Skill Bloom Skill
Question Num. Words Question Num. Words Question Num. Words Question Num. Words
Answer Num. Chars Answer Num. Chars Answer Num. Chars Answer Num. Chars
Num. Reasoning Steps Num. Reasoning Steps Num. Reasoning Steps Num. Reasoning Steps
MDL MDL MDL MDL
Table 1: Hardness measures we use for each dataset. Grayed-out options are not present in the dataset annotations, and thus not used in our experiments.
3 Measuring Datapoint Hardness
Measuring easy-to-hard generalization requires drawing a distinction between easy and hard data, defined in terms of human ability to correctly label the data. There could be many ways to describe what makes problems harder, including that (1) only people with specialized training and knowledge can solve the problem; (2) it takes people longer to solve the problem; (3) people are less certain that their final solution is correct; (4) people with similar expertise naturally disagree about the solution to the problem, while agreeing that there is an objective solution; (5) experts know of a reliable method for obtaining the answer to a problem, but it is costly in terms of time and effort or possibly noisy in its outputs (like conducting scientific experiments). Each of these dimensions clearly reflects that hardness lies on a spectrum.
In this paper, we aim to capture the above properties in a number of specific measures we can obtain for each instance in our datasets, including:
  1. 1.
    Education/Grade Level: What level of education (possibly in a particular domain) would typically lead one to be able to answer the question?
  2. 2.
    Expert Rating: How difficult would an expert rate the question, on an ordinal scale?
  3. 3.
    Required Cognitive Skill: What cognitive skills are required to answer the question? This rating is based on Bloom’s taxonomy of cognitive skills, which are ranked in order of increasing complexity: (1) Remembering, (2) Understanding, (3) Applying, (4) Analyzing, and (5) Evaluating Bloom et al. (1956); Adams (2015).
  4. 4.
    Question Num. Words: Question length is a natural proxy for question hardness, as longer questions can involve more premises or a greater number of concepts.
  5. 5.
    Answer Num. Chars: We also consider Answer Num. Chars, since longer answers may reflect more specific or more complex problems. Character count provides a measure that is applicable across tasks.
  6. 6.
    Compositional Steps: Compositional reasoning is more difficult than executing individual reasoning “primitives,” since it requires first getting the primitives correct and then combining their solutions correctly. We consider how many individual reasoning steps are involved in answering a question (i.e., the number of subproblems whose solutions must be combined), according to human-annotated reasoning chains.
These notions of hardness are fundamentally human notions of hardness, but we can also measure a model-based metric for datapoint hardness. In this direction, we use a minimum-description-length (MDL) metric Voita and Titov (2020). In practice, MDL can be measured by computing a test datapoint’s average label probability across models of identical architecture trained on increasing quantities of training data for a task Perez et al. (2021). Intuitively, MDL captures how hard on average an in-distribution test datapoint is for a model to generalize to given some amount of training data. Ultimately, we use our MDL metric to capture how well a stronger model generalizes to data that is hard according to a weaker model, in order to simulate a setting where humans cannot label hard problems that they would like for a strong model to solve. Details for MDL computation are given in Appendix A.
  1. 7.
    Minimum Description Length: A model-based measure of datapoint hardness, given by the datapoint’s label probability averaged over different models trained with increasing amounts of training data. Datapoints with lower average label probabilities are considered harder.
In experiments, we use four datasets that contain instance-level annotations for some portion of these measures, as shown in Table 1.
  1. $\bullet$
    ARC Clark et al. (2018): U.S. gradeschool science questions in multiple-choice format. We combine ARC-Easy and ARC-Challenge splits. Random performance is 25%.
  2. $\bullet$
    MMLU Hendrycks et al. (2020): Domain-specific multiple-choice questions for a large number of domains. We subset to high school and college level math, physics, biology, chemistry, and computer science questions (termed MMLU-STEM-5). Here, grade level is high school (HS) vs. college. See Figure 2 (left) for an example. Random performance is 25%.
  3. $\bullet$
    StrategyQA Geva et al. (2021): Yes/no general knowledge trivia questions requiring compositional reasoning over individual facts. The “Num. Reasoning Steps” measure is the number of facts that must be combined. Majority-class vote performance is 53.9%.
  4. $\bullet$
    GSM8k Cobbe et al. (2021): U.S. grade school math word problems in direct answer format (i.e., no answer choices given). Random performance is 0%. The number of steps in the solution to the problem forms the “Num. Reasoning Steps” measure and is obtained from the human-annotated reasoning chain collected for each problem. See Figure 2 (right) for an example.
There are generally fewer hard datapoints than easy datapoints in our datasets, given the relative difficulty of collecting hard data. In MMLU-STEM-5, for example, there are 603 college level questions and 1143 high school questions. We show histograms for each hardness measure distribution in Appendix Fig. 20 (see Appendix C for further dataset information).
4 Experiment Setup
Models. We conduct experiments with Llama-2 base models, for sizes 7b, 13b, and 70b Touvron et al. (2023b). In Appendix B, we show results for an RLHF version of Llama-2-70b, Qwen-72b Bai et al. (2023), and Mixtral-7x8b Jiang et al. (2024). Results are robust across models.
Data Hardness Stratification. To separate datasets into easy and hard data (with leftover data being medium data), we define easy/hard cutoffs as follows: for Question Num. Words, Answer Num. Chars, and MDL, we automatically define these values to be at the 30th and 70th percentiles of the variable range. Other variable cutoffs are defined manually: For ARC, Grade Level is easy (3-5), medium (6-7), hard (8); Difficulty Score is easy (1), medium (2), hard (3); Bloom Skill is easy (1-2), medium (3), hard (4-5). For MMLU, Grade Level is easy (high school) and hard (college) with no medium. For StrategyQA, Num. Reasoning Steps is easy (1-2), medium (3), hard (4-5). For GSM8k, Num. Reasoning Steps is easy (2-3), medium (4-5), hard (6-11).
Training Methods. We fit models to data with in-context learning (ICL; Brown et al., 2020), linear classifiers trained on frozen model hidden states Belinkov (2022), or QLoRA Dettmers et al. (2023). StrategyQA and GSM8k benefit heavily from utilizing chain-of-thought reasoning (CoT; Wei et al., 2022), so we primarily conduct experiments for these datasets with ICL+CoT and QLoRA+CoT (using reasoning chains from the datasets). See descriptions of each method below, with full detail in Appendix D.
Figure 3: Correlation heatmaps for ARC and GSM8k hardness measures.
Figure 4: Accuracy vs test data hardness across datasets (using Llama-2-70b with ICL). Data that humans find harder is also harder for LMs.
  1. 1.
    ICL: We leverage the in-context learning ability of LMs by placing “training” data into a prompt preceding each test question. We use $k=10$ in-context examples for ARC and MMLU and $k=8$ examples for StrategyQA and GSM8k (we see diminishing returns for larger $k$). When scoring multiple choice questions (no CoT), we get a model prediction by computing the answer probability for each answer choice given the test input and the prompt. When generating outputs with CoT, we greedily generate up to $t=100$ tokens for StrategyQA and $t=300$ tokens for GSM8k, then parse the string to get the predicted answer and compute exact-match accuracy against the label.
  2. 2.
    Linear Probing: We train a linear classifier on frozen LM hidden states. This is an effective method for performing multiple choice QA using LM representations Liu et al. (2023), and it does not require any finetuning of the underlying LM. For a given question, we compute one representation per answer choice by concatenating the question and answer choice and feeding it to the model. Then, we score each representation $z$ by applying the linear probe: $f(z;w)=w^{T}z$. The answer choice with the highest score is returned as the prediction. The probe weight $w$ is trained using SGD to minimize cross-entropy loss on a dataset of representations $Z=\{\{z_{i,j}\}_{j=1}^{j=|A|}\}_{i=1}^{N}$ derived from $N$ training datapoints with $|A|$ answer choices. Note we are not able to use this method for GSM8k as there are not multiple answer choices per datapoint.
  3. 3.
    QLoRA: Recently, low-rank adaptation of specific model parameters has enabled extremely parameter-efficient finetuning of LMs. To finetune our LMs, we execute QLoRA with the LoRA implementation from HuggingFace peft Mangrulkar et al. (2022) and the 8-bit AdamW from bitsandbytes Dettmers et al. (2022). We train the default layers for Llama-2 with rank $r=16$ adapters, $\alpha=32$, and dropout $p=0.1$. Model predictions are obtained in the same manner as for ICL, i.e., by scoring multiple choice options or generating $t=100/300$ tokens for StrategyQA/GSM8k.
Unsupervised Baseline. Our unsupervised baseline is zero-shot prompting, scoring the answer choice probabilities given the question and taking the highest probability answer as the model prediction. The one exception to this is for GSM8k, which does not have multiple answer choices per question. For this dataset, we use a simple “Let’s think step by step” style prompt. See Appendix D.
Training Size Controls. For all experiments with linear probing and QLoRA, we use $n=160$ train points. While we would prefer to use more training data, the bottleneck we face is that fairly comparing easy-to-hard with hard-to-hard generalization requires both fixing the amount of training data and leaving enough hard data left over for testing. Since we have as few as $n=603$ hard test points for MMLU, we have to limit training data to $n=160$ points to leave enough test data for reasonably small confidence intervals. Linear probing and QLoRA demonstrate good sample-efficiency when applied to Llama-2-70b, so we are able to obtain comparable (and sometimes better) performance than ICL across datasets using these methods.
Statistical Testing. We perform experiments using 5 random seeds, controlling the training data selection (leaving remaining data for testing). To obtain confidence intervals and compute $p$-values for comparing results, we perform a block bootstrap to combine test and train data variance Efron and Tibshirani (1994). See Appendix E for further detail.
5 Experiments
Figure 5: Accuracy on hard test data vs train data hardness across datasets (using Llama-2-70b and ICL). Models recover 70-100% of the supervision gap (between Unsupervised and Hard) by training on Easy data.
5.1 RQ1: How Can We Measure Data Hardness? Do Different Approaches Agree?
We first explore the properties of our hardness measures for each dataset, to understand how they stand in comparison to one another.
Design. We first create correlation heatmaps for the hardness measures in our datasets, using a Spearman rank-order correlation Spearman (1987) between hardness values for each datapoint in a dataset. We show examples for ARC and GSM8k, which together provide complete coverage of our seven notions of data hardness (results for all datasets and hardness measures shown in Appendix Fig. 10). To validate that these measures are meaningful, we assess model performance across test data hardness levels, using Llama-2-70b and ICL with randomly sampled prompt examples.
Results. Based on Fig. 3, we find that our hardness measures are empirically very diverse. Off-diagonal correlations are usually fairly low, suggesting that these measures capture different possible aspects of datapoint hardness. [^3] Some of the ARC hardness correlations are low by design, as grade school exam questions are designed to vary in hardness within grade level (note the correlation of $\sim$0 between “Expert 1/2/3 Difficulty” and Grade Level). However, it appears that all of our hardness measures meaningfully capture some aspect of datapoint hardness, as model accuracy declines for harder test data for each of these measures (shown in Fig. 4), including for model-based hardness as measured by 7b-parameter models (shown in Appendix Fig. 11). Interestingly, we find that our model-based hardness measure is quite distinctive from our human-based measures, reaching a correlation with human hardness of at most 0.14 for ARC. This suggests that it is important to avoid relying only on model-based measures of hardness. Overall, we conclude that easy-to-hard generalization should be assessed with multiple notions of human hardness that are informed by cognitive science, since it appears that several meaningful measures of “hardness” do not strongly correlate with one another.
5.2 RQ2: Can We Do Well on Hard Data by Training on Easy Data?
We now examine how well models generalize from easy training data to hard test data.
Design. For each of our hardness measures, we test models on exclusively hard test data (according to that hardness measure), while varying whether they are trained on easy or hard data. We show results for Llama-2-70b using ICL. For results with additional hardness measures, see Appendix Fig. 19, and for other training methods, see Appendix Fig. 15. We focus on hard test data here, since that is the setting we are most interested in, but we provide results testing on the full data distribution in Appendix Fig. 13 and easy test data in Fig. 14.
Results. Surprisingly, Llama-2-70b shows comparable generalization to hard test data regardless of whether it is trained on easy or hard data (Fig. 5). In fact, across all six hardness measures, the Supervision Gap Recovered is between 70% and 100% (see SGR values with CIs and statistical significance tests in Appendix Table 2). Interestingly, for ARC and MMLU, there is no difference in easy vs. hard generalization using ICL. These results are robust across training methods (see Appendix Fig. 15). With QLoRA, the SGR also remains within 70%-100% for ARC, MMLU and StrategyQA. While GSM8k appears to exhibit worse easy-to-hard generalization, we note that easy-to-all generalization is actually equally good to hard-to-all generalization (see Fig. 13). Thus it seems like easy data provides surprisingly good supervision for LMs.
These results contrast notably with past work in curriculum learning and compositional generalization Bengio et al. (2009); Lake and Baroni (2018). This is likely because the models we consider, like Llama-2-70b, have learned much more during pretraining than models commonly used in work on curriculum learning and compositional generalization. So, it would seem that training these models on relatively small amounts of easy data successfully elicits the relevant task knowledge from the models in a way that is largely invariant to datapoint hardness.
Figure 6: Hard test accuracy is often higher when training on comparable amounts (costs) of easy rather than hard data. Here, we suppose hard data is 2x costlier to collect. Results are for Llama-2-70b with linear probes.
5.3 RQ3: What Are the Cost-Benefit Tradeoffs of Collecting Easy vs. Hard Training Data?
One implication of the results from Sec. 5.2 is that if easy data is almost as good as hard data, it could be better to collect and train on easy data, since hard data can be noisier and costlier to collect Sambasivan et al. (2021). Hence, we test the hypothesis that training on easy data outperforms hard data under two possible assumptions: (1) that one can collect more easy data than hard data given a fixed budget (i.e., time, money), and (2) that easy data is less noisily labeled than hard data.
Design. For (1) the budget assumption, we train linear probes on either easy or hard data using datasets of sizes in {10, 20, 40, 80, 160, 320}. We then show hard data test performance vs. training cost, assuming hard data costs twice as much as easy data to collect, meaning labeling 40 easy training points is equivalent in cost to labeling 20 hard training points. For (2) the noise assumption, we assume that easy data is mislabeled $p\%$ of the time, while hard data is mislabeled $2p\%$ of the time. Here, we measure test performance on hard data given different values of $p$. Note the 1:2 data collection cost ratio is almost exactly the ratio observed in MMLU-STEM-5, which contains 603 college level questions and 1143 high school questions, and a 1:2 labeling error ratio is plausible as well given expert human accuracy on datasets like MMLU (estimated at 89.8%) and GPQA (estimated at $\leq$72for difficult graduate level STEM questions) (Hendrycks et al., 2020; Rein et al., 2023). We use linear probes for this experiment given the expense of running 5 seeds of CoT experiments across different QLoRA training sizes for StrategyQA.
Figure 7: When hard train data has noisier labels than easy train data, training on easy data can give better hard test performance (shown for MMLU-STEM-5 using Llama-2-70b with a linear probe).
Figure 8: Models perform as well on hard MMLU data when prompted with easy MMLU data as they do when prompted with hard data, regardless of model size ($k=10$ examples used for ICL).
Results. We show results for the data budget assumption in Fig. 6. There is more often than not a clear advantage to fitting a model to easy data over hard data when the cost ratio between them is 1:2. The exceptions are for Bloom Skill and StrategyQA reasoning steps, where it appears that hard data does eventually meet or surpass the efficacy of easy training data. In the other three settings however, one can improve the model accuracy on hard test data by using easy rather than hard training data.
For the noise assumption, we draw a similar conclusion based on the results for MMLU in Fig. 7. It can be better to train on easy data rather than hard data, even when testing on hard test data. In this case, easy data becomes preferable when its labeling error rate is 10%, meaning the error rate for hard data is 20%. While these error rates are somewhat large, they are quite possible for difficult domain questions (Rein et al., 2023), suggesting that there are plausible settings where it is better to train on easy data than hard data due to label noise.
5.4 RQ4: Is Easy-To-Hard Generalization Consistent Across Model Scale and Train-Test Hardness Gap Size?
We attempt to answer two questions likely to be relevant as models become more capable: (1) how does the Supervision Gap Recovered change as models scale up, and (2) how does hard test performance change as the gap in train and test hardness grows? We are interested in these questions because in various settings AI performance may exceed human expert performance, and we want to know whether it will become more and more difficult to supervise models as this occurs.
Design. For question (1), we show easy-to-hard and hard-to-hard performance on MMLU for three different model sizes of Llama-2, namely 7b, 13b, and 70b. For question (2), we test models on hard MMLU data (college STEM questions), while training them on: high school MMLU data, 8th grade ARC data, 3rd grade ARC data, and no training data (unsupervised).
Results. First, we find that models show similar levels of Supervision Gap Recovered across scale (Fig. 8). SGR is 109% for the 7b model, 75% for the 13b, and 97% for the 70b, while the absolute difference between easy-to-hard and hard-to-hard performance is never more than 0.4 points. For our second question, we find that the difference between train and test hardness may have an effect on test performance (Fig. 9). Across methods, we see some decline in generalization once the gap between train and test hardness becomes sufficiently large, however 3rd grade supervision can be surprisingly effective for college STEM questions (e.g. SGR falls from 74% to 57% when training on 8th grade vs. 3rd grade questions using ICL). In an evaluation across our other hardness measures, easy training data is only marginally worse than medium training data (see Appendix Fig. 17). Together these results suggest that, while (1) the scalable oversight problem does not become harder as models scale up, (2) easy-to-hard performance may begin to decline when the gap between train and test hardness becomes sufficiently large.
Figure 9: Performance on hard test data may begin to decline when training on very easy data. Results shown for college-level STEM questions using Llama-2-70b (see other hardness measures in Appendix Fig. 17).
6 Discussion
Are Our Tasks Hard Enough to Provide Generalizable Easy-To-Hard Results? While humans can and have effectively labeled all of the datasets we use, the authors of this paper certainly could not have. Already, benchmark datasets for LMs require specialized domain expertise as benchmarks have increased in difficulty over time Clark et al. (2018); Hendrycks et al. (2020). The largest difficulty gap that we test in this paper is between 3rd grade and college level STEM questions (using ARC and MMLU). Concurrent work has called for work studying gaps as large as 3rd grade to 12th grade (described as “huge leaps in generalization”; Burns et al., 2023), and MMLU has been used in the past for scalable oversight experiments Bowman et al. (2022). Therefore, we see our results as relevant for future work that may operationalize “easy” and “hard” differently. One other interesting consideration here is that label noise in increasingly hard datasets may make experimental results less reliable as we test on harder and harder test sets, which could be an argument in favor of conducting experiments specifically on difficulty ranges like in MMLU where we have high trust that college-level test questions are labeled correctly.
How Do LMs Solve Hard Problems From As Few As Ten Easy Examples? Language models are known to be highly sample-efficient learners Brown et al. (2020), but our results demonstrate that they also efficiently learn to solve hard problems from easy data (recall that ICL achieves SGR scores of 70%-100% using no more than $k=10$ examples). How is this degree of supervision sufficient for recovering hard-to-hard performance? Evidently, training on even small amounts of easy data successfully elicits relevant knowledge from LMs in a way that is largely invariant to datapoint hardness. This could be because this kind of training encourages models to answer questions based on “truthfulness” representations of text, which should be invariant across domain and data hardness (see Azaria and Mitchell, 2023; Levinstein and Herrmann, 2023; Marks and Tegmark, 2023). We would point out that we do not interpret our results as models merely “learning the task format” or “learning the input/output space” as opposed to true generalization, because we are able to generalize to MMLU-STEM-5 college questions (a diverse mixture of physics, chemistry, biology, mathematics, and computer science questions) by using 3rd grade or 8th grade questions from ARC (Fig. 1), which come from a different dataset entirely. Hence it appears that training on easy data encourages models to give correct outputs for hard questions, though we note that the training signal is still underspecified in cases where SGR is less than 100%.
Why Not Use Test Questions About Knowledge From Beyond Pretraining Data Cutoffs? We want to emphasize that we are not trying to train models to answer questions for which there is no relevant information in their (pre)training data. Rather, we are trying to elicit knowledge from models that we suspect they may know, using fundamentally weak supervision. This is why hard-to-hard performance provides an upper bound for easy-to-hard generalization in our experiments (if a model simply cannot answer a question, the upper bound should reflect that). It would be interesting to consider test performance on “future” data given training data from before some cutoff, such that “easy” data is from the past and the “hard” training data is relevant “future” data.
7 Conclusion
We study the problem of easy-to-hard generalization, showing that (1) several meaningful human and model-based hardness measures disagree about which data is hardest; (2) LMs trained on easy data often perform nearly as well as those trained on hard data, recovering 70-100% of the Supervision Gap between an unsupervised lower bound and hard-to-hard upper bound; (3) practically, one can perform better on hard test data by collecting and training on easy data rather than hard data when the hard data is noisier or costlier to collect; and (4) SGR may begin to decline when the gap between train and test hardness becomes sufficiently large. These results are robust across datasets, training methods, hardness measures, and model size. Our findings suggest that the scalable oversight problem may be easier than previously thought.
Acknowledgements
We thank Yanai Elazar, Kyle Richardson, Ben Bogin, and Tushar Khot for helpful feedback on this paper, as well as Joris Baan, Nouha Dziri, and others at the Allen Institute for AI for useful conversations about the work. We especially thank Oyvind Tafjord for assistance regarding the human hardness measures we use. This work was conducted while Peter Hase was a research intern at the Allen Institute for AI.
Appendix A Measuring Minimum Description Length
In addition to our human hardness measures, we employ a model-based metric based on minimum-description-length Voita and Titov (2020); Swayamdipta et al. (2020); Perez et al. (2021). Since experiments use up to 70b parameter LMs, we measure MDL with models in the 7b parameter range, including Falcon-7b TII (2023), Mistral-7b Mistral AI (2023), Persimmon-8b Elsen et al. (2023), and Llama-1-7b Touvron et al. (2023a). To get one MDL per datapoint, we average the MDL scores obtained for each of the four models. For a single model, we obtain a score by training $N$ models on training sizes in $n\in\{5,20,80,340,900\}$ (roughly log-uniform) when fitting models with linear classifier heads or QLoRA and averaging model label confidences across these $n$ per-datapoint scores. For ICL we compute MDL using no training data, i.e. $n=0$. In this way, our Linear Probe and QLoRA MDL metrics represent MDL according to the theoretical definition which involves increasing amounts of training data, while MDL (ZS Prompt) represents the confidence that 7b models assign to labels for data with no supervision. All metrics are then used for assessing how stronger models will perform on data that weaker models find to be hard. See distributions of MDL scores on each dataset in Fig. 20. We only measure probing-based and QLoRA-based MDL for ARC and GSM8k, where we have sufficient data to set aside $n=1000$ points (up to 900 for training and 100 for model validation).
Appendix B Additional Results
We include a number of additional results in this section.
  1. 1.
    For a table of model accuracies by training method in an all-to-all setup, see Table 4.
  2. 2.
    We show correlations between hardness measures for all data in Fig. 10.
  3. 3.
    We show that test accuracy declines with test data hardness for QLoRA in Fig. 12, and using ICL with additional hardness measures in Fig. 11.
  4. 4.
    We show easy-to-hard generalization as measured on all test data (not subsetting to hard test data) in Fig. 13, as well as testing on easy test data in Fig. 14.
  5. 5.
    We show SGR statistics for all training methods in Fig. 15.
  6. 6.
    We give SGR estimates along with confidence intervals and $p$-values, obtained by block bootstrap, in Table 2.
  7. 7.
    We show easy-to-hard generalization on StrategyQA across different models in Fig. 16.
  8. 8.
    We give additional results for test performance by multiple training difficulty levels in Fig. 17.
  9. 9.
    We give test performance by training difficulty for multiple $\sim$70b models in Fig. 18.
Figure 10: Correlations between hardness measures across datasets. We omit MDL for GSM8k because 7b parameter models obtain extremely high loss on GSM8k problems, and MDL is valid as a metric only when using a reasonably good model of the data.
Figure 11: Test accuracy declines as test data hardness increases (shown for additional hardness measures).
Figure 12: Test accuracy declines as test data hardness increases (shown for QLoRA with Llama-2-70b).
Figure 13: Easy-to-hard generalization measured on all test data (not subsetting to hard test data), while training on easy/hard data defined according to each hardness measure. Results are similar to testing on hard data, except for GSM8k, where accuracy on the whole data distribution becomes comparable (training on easy data outperforms hard data on easy/medium test data).
Figure 14: Easy-to-hard generalization measured on easy test data, while training on easy/hard data defined according to each hardness measure. This plot shows hard-to-easy generalization for each dataset, compared to easy-to-easy genearlization. On some datasets, hard data makes for better training data, while for others, easy training data is better for easy test performance.
Figure 15: The Supervision Gap Recovered (SGR) shown by training method. Easy-to-hard generalization varies somewhat by training method used, but SGR remains surprisingly high across datasets for the two most effective training methods, ICL and QLoRA.
Figure 16: Easy-to-hard generalization results for different models on StrategyQA. Results are similar for Llama-2-70b and Mixtral, while Qwen appears to have been trained on StrategyQA data in a CoT format.
Figure 17: Test performance on hard data sometimes declines more significantly as the gap between train and test hardness grows, but often the difference between training on Medium and Easy data is relatively small in nature. MMLU not shown here since there are only two hardness levels for that dataset (high school vs. college). See Fig. 9 for more results training on college vs. high school vs. 8th grade vs. 3rd grade data.
Figure 18: Test performance on hard data declines as the gap between train and test hardness grows for reasoning datasets, across models, using ICL with $k=10$.
Figure 19: Easy-to-hard generalization for additional hardness measures for each dataset, using Llama-2-70b with ICL. SGR values remain high across possible hardness measures, meaning easy data provides surprisingly good supervision. We do not represent Answer Num Chars. for StrategyQA here because that would cleanly divide the data into ‘no’ and ‘yes’ categories. We do not conduct any additional experiments for GSM8k hardness measures as these experiments (involving sampling CoTs with $t=300$ tokens) are extremely computationally expensive.
Dataname Hardness Measure SGR Estimate Test Hardness $n$
ARC Grade Level 0.96 ± 0.10 ($p$ < 1e-4) Hard 1588
ARC 1/2/3 Difficulty 0.98 ± 0.36 ($p$ = 0.0033) Hard 1588
ARC Bloom Skill 1.00 ± 0.20 ($p$ < 1e-4) Hard 1588
MMLU HS vs. College 0.97 ± 0.59 ($p$ = 0.0158) Hard 603
StrategyQA Num Reasoning Steps 0.72 ± 0.93 ($p$ = 0.0788) Hard 427
GSM8k Num Reasoning Steps 0.79 ± 0.60 ($p$ = 0.0125) Hard 333
ARC Grade Level 1.00 ± 0.09 ($p$ < 1e-4) All 3521
ARC 1/2/3 Difficulty 0.96 ± 0.08 ($p$ < 1e-4) All 3521
ARC Bloom Skill 0.98 ± 0.08 ($p$ < 1e-4) All 3521
MMLU HS vs. College 1.00 ± 0.27 ($p$ = 0.0001) All 1746
StrategyQA Num Reasoning Steps 0.87 ± 0.32 ($p$ < 1e-4) All 2290
GSM8k Num Reasoning Steps 0.98 ± 0.39 ($p$ = 0.0003) All 2065
Table 2: Supervision Gap Recovered (SGR) statistics for Llama-2-70b with ICL, on hard test data or all test data, defined per dataset and hardness measure. Confidence intervals are 95% CIs estimated by block bootstrap (accounting for test data variance and train data variance), and $p$-values represent a test for a difference from 0.
Appendix C Dataset Details
We provide additional details for each dataset below, including test data subsetting decisions (see final sample sizes in Table 2). All datasets are publicly available. See also Table 1 for a list of which hardness measures are available for which datasets, as well as sample sizes. See Fig. 20 for histograms of hardness measures per dataset.
  1. $\bullet$
    ARC Clark et al. (2018): [^4] https://huggingface.co/datasets/ai2_arc U.S. gradeschool science questions in multiple-choice format. We combine the ARC-Easy and ARC-Challenge splits. We release metadata including human hardness metadata accompanying the original source of the questions in our codebase at https://github.com/allenai/easy-to-hard-generalization. We set aside 1000 points for MDL computation, so experiments are conducted on $n=3521$ test points.
  2. $\bullet$
    MMLU Hendrycks et al. (2020): [^5] https://huggingface.co/datasets/tasksource/mmlu Domain-specific multiple-choice questions for a large number of domains. We subset to high school and college level math, physics, biology, chemistry, and computer science questions (termed MMLU-STEM-5). Here, grade level is high school (HS) vs. college. See Figure 2 (left) for an example.
  3. $\bullet$
    StrategyQA Geva et al. (2021): [^6] https://huggingface.co/datasets/wics/strategy-qa General knowledge trivia questions requiring compositional reasoning over individual facts. The number of facts that must be combined forms the “Num. Reasoning Steps” measure.
  4. $\bullet$
    GSM8k Cobbe et al. (2021): [^7] https://huggingface.co/datasets/gsm8k U.S. grade school math word problems. The number of steps in the solution to the problem forms the “Num. Reasoning Steps” measure and is obtained from the human-annotated reasoning chain collected for each problem. We set aside 1000 points for MDL computation, then further subset to about $n=2000$ test points given the extreme expense of sampling CoTs with $t=300$ tokens for 70b parameter models. See Figure 2 (right) for an example datapoint.
Figure 20: Distributions for hardness measures for each dataset and hardness measure.
Distributions for hardness measures for each dataset and hardness measure (from Table. 1) are shown in Fig. 20.
Appendix D Modeling and Tuning Details
We provide additional information around model tuning for each training method here. All training methods use the same prompts, one per dataset, that were selected based on their success in past work Saha et al. (2023). We show prompts for ARC and MMLU in Table 3. In this template, the {} placeholders are filled with the question, four answer choices, and a single answer choice (explained next). We use this prompt for multiple choice scoring of the four answer choices for all methods, meaning that we run four forward passes to either (1) compute answer choice probabilities for each answer choice for ICL and QLoRA, or (2) collect final answer choice token representations for each answer choice for linear probing. Thus the final answer slot, “A: {}” is populated by each answer choice once. When prompting with $k$ in-context examples for ICL, we separate examples with a line break (one line between each pair of examples).
For StrategyQA and GSM8k, we use a different prompt format for CoT, shown in Table 3. In this template, the {} placeholders are filled with the question, the human reasoning chain, and the answer choice (only for in-context examples). This prompt is used to generate new reasoning chains and answers at test time, so there is no text included after “A:” for the test input. When prompting with $k$ in-context examples for ICL, we separate examples with a line break (one line between each pair of examples). The exception to this formatting is for GSM8k’s Unsupervised Baseline, which uses a “Let’s think step-by-step” prompt (we also considered this for StrategyQA, but zero-shot answer choice scoring worked better). The step-by-step prompt is shown in Table 3. The test input is supplied to the curly brace placeholder.
ARC+MMLU prompt
Question: {}
A) {}
B) {}
C) {}
D) {}
Answer: {}
StrategyQA+GSM8k prompt
Q: {}
A: {} So the answer is {}
Unsupervised GSM8k prompt
Q: [question text here]
A: Let’s think step by step.
1. [step one]
2. [step two]
N. [last step]
Therefore, the answer is [answer here].
Now you try!
Q: {}
A: Let’s think step by step.
1.
Table 3: Prompt formats used in this paper. Question text, reasoning text, answer choices, and answer text are imputed in curly brackets. The notation “[step one]” is literal, and no variables are imputed in these brackets. When in-context examples are included in the prompt, we separate each example with one empty line.
For ICL, we select $k=10$ for ARC and MMLU and $k=8$ for StrategyQA and GSM8k as we see diminishing returns to accuracy from larger values of $k$, and using larger $k$ values significantly slows down experiments.
For Linear Probing, we fit a linear classifier to frozen LM hidden states. For a given question, we compute one representation per answer choice by concatenating the question and answer choice and feeding it to the model. To get a single representation from the LM forward pass, we concatenate the representations at the last token index (i.e., the last answer token) from the middlemost and last layer. Then, we score each question-answer choice representation $z$ by applying the linear probe: $f(z;w)=w^{T}z$. The answer choice with the highest score is returned as the prediction. The probe weight $w$ is trained using SGD to minimize cross-entropy loss on a dataset of frozen representations $Z=\{\{z_{i,j}\}_{j=1}^{j=|A|}\}_{i=1}^{N}$ for a dataset of $N$ training datapoints and $|A|$ answer choices. We optimize the weight with the default SGD implementation in PyTorch Paszke et al. (2019) for 100 epochs, without early stopping on any dev data. Based on tuning experiments with Llama-2-13b on ARC, we chose to use SGD rather than AdamW, selecting the number of epochs for convergence, and we chose to use the middlemost and last layer representations (concatenated) rather than either on its own. Note this produces very high dimensional inputs, but by using $\ell_{2}$ regularization with $\lambda=1$, we are able to stably fit probes to these $2d$-dimensional input representations (where $d=8192$ for Llama-2-70b) with as few as $n=10$ training points. The learning rate was fixed at 5e-2.
For QLoRA, we selected hyperparameters based on early experiments with Llama-13b on ARC. Based on this setting, we selected an adaptor rank of $r=16$ rather than $r=8$, with default hyperparameters otherwise, including default selected layers to optimize. The learning rate was fixed at 1e-4, and the model is optimized with a batch size of 50. At least 10 gradient updates are performed, or a minimum of three epochs, whichever yields more gradient updates. This means that for train $n=160$ points and a batch size of 50, we generally perform 12 gradient update steps (3 epochs) in our experiments.
To select the number of decoding steps for each datasets ($t=100$ for StrategyQA and $t=300$ for GSM8k), we wanted to make sure that we were generating reasoning chains long enough for Llama-2-70b to solve hard test questions. Therefore, we intentionally selected this parameter based on hard test performance, in order to use as small a valuable as possible (based on experiment efficiency) that did not compromise test performance on hard data.
All models are run in 8bit quantization, except for Qwen and Mixtral, which are run in 16bit format, and falcon-7b and persimmon-8b, which are run in full precision. We observe no performance loss from quantization in any experiment.
Dataset Method CoT $n$ Acc (%)
ARC ICL No 0 85.94
ARC ICL No 10 91.77
ARC Linear Probe No 160 89.48
ARC QLoRA No 160 89.47
GSM8k ICL Yes 0 4.80
GSM8k ICL Yes 8 56.24
GSM8k QLoRA Yes 160 52.64
MMLU-STEM-5 ICL No 0 48.30
MMLU-STEM-5 ICL No 10 56.83
MMLU-STEM-5 Linear Probe No 160 53.01
MMLU-STEM-5 QLoRA No 160 52.77
StrategyQA ICL No 0 62.40
StrategyQA ICL No 4 68.89
StrategyQA ICL Yes 8 72.86
StrategyQA Linear Probe No 160 68.79
StrategyQA QLoRA No 160 66.36
StrategyQA QLoRA Yes 160 75.09
Table 4: Model accuracy table when trained on randomly selected data from the whole data distribution and tested on the whole data distribution (zero-shot ICL rows included), using Llama-2-70b. Averaged over 5 seeds. Compare to Fig. 13.
Appendix E Statistical Testing
Here, we describe in greater detail how our statistical testing works. We aim to make the most of the data we have, e.g. $n=603$ hard datapoints for MMLU-STEM-5. Ultimately, we want to use a block bootstrap that resamples (1) test datapoints and (2) models (equivalent to picking which random seed was chosen for training), in order to account for variance due to limited test data, selection of training data, and any random training dynamics. Therefore, we run five random seeds for each experiment, randomly selecting training data and using all remaining data as test data. Each experiment produces a matrix (block) of results, with up to five model predictions per datapoint. Running this matrix through a block bootstrap that resamples rows and columns produces a confidence interval for the statistic of interest Efron and Tibshirani (1994). When computing the mean of a resampled matrix, we ignore missing values (which represent that a datapoint was used for training and could not be tested on). We take 100,000 resamples.
We can use a bootstrap to obtain estimates for our SGR statistic too. We aim to estimate the expected value of the random variable
$\vspace{-1pt}\frac{\textrm{Easy}-\textrm{Unsupervised}}{\textrm{Hard}-\textrm{Unsupervised}}\vspace{-1pt}$
using the samples Unsupervised, Easy, and Hard representing their respective experiment outputs ($n\times 5$ matrices of model predictions). We perform a paired test with respect to test datapoints (resampling the same rows across each matrix), while not assuming random seeds are paired (resampling different columns for each matrix). Note that for the Unsupervised matrix, each column is identical because there is no prompt data. The results of this analysis are given in Table 2. When showing results for hard test data, we subset to just points that are hard according to their respective hardness measure.